In case of headache, 3 patients out of 5 take Panadol while 2 out 5 take another medication M. 75% of patients feel better knowing that they took Panadol. 90% feel better given they took the other medication


1. What is the probability of having no longer a headache? 2. What is the probability that a patient feels better given he took Panadol?


1
Expert's answer
2020-12-21T19:01:00-0500

Let "A" be the event "patient takes Panadol", "B" be the event "patient takes another medication M", and "C" be the event "patient feels better".

Given that

"P(A)=\dfrac{2}{5}=0.4, P(B)=\dfrac{3}{5}=0.6,"

"P(C|A)=0.75, P(C|B)=0.90"

1. By the theorem of total probability 


"P(C)=P(C|A)P(A)+P(C|B)P(B)"

"=0.75(0.4)+0.90(0.6)=0.84"

The probability of having no longer a headache is "0.84."


2.


"P(C|A)=0.75"

The probability that a patient feels better given he took Panadol is "0.75."